OpenStudy (anonymous):

What is the area of a triangle with a base of 3.2 cm and a height of 9.6 cm? A. 6.4 cm² B. 12.8 cm² C. 15.36 cm² D. 30.72 cm²

OpenStudy (_michaeld22):

A = 1/2Bh. 3.2*9.6 = 30.72 30.72/2 = ?

OpenStudy (dannyrod2000):

C

OpenStudy (dannyrod2000):

you could use A=1/2*Bh like @_Michaeld22 gave you